HMV 904 #2 Restoration
Cabinet. (pictures) The cabinet is in fair condition. We will have it refinished.
Chassis. (pictures) It has a single chassis, which is in good condition. It has a single chassis, which is in good condition. This chassis is much more complicated and crowded than most pre-1945 sets. We will carefully clean the chassis with water and detergent, then use mineral spirits to remove stubborn grease.
All paper capacitors will be rebuilt (see the procedure for this). Each electrolytic capacitor will be tested for leakage and capacity. If bad, new electrolytics will be installed inside the old ones. The high voltage (EHT) capacitor also has to be rebuilt.
The high voltage (EHT) transformer started overheating as we were testing the power supplies. A new one was fabricated by Ed Dinning. It is almost exactly the same physical dimensions, but in order to make it look more like the original, we removed some parts from the original and put them on the new one. Here is a detailed description.
All of the capacitors have been rebuilt, and we started working our way through the set. The first problem was that the black spray paint used on the high voltage (EHT) transformer was conducting, and causing arcing. After cleaning it off the high voltage terminals the problem was solved.
The next step was to install the CRT and check the deflection circuits. They worked, but there was insufficient height and width. We decided to come back to that problem later.
Then, we set the local oscillator to 37 mHz, which required adding a small (5 pf) capacitor across the oscillator coil. We then had video at the CRT cathode. The sync was not working properly, but we could see an image on the CRT. After about ten minutes we noticed that the video level had dropped significantly. A check of the first and second IF tube (valve) voltages revealed that the plate and screen grid voltages were well below normal, and the cathode voltages were well above normal. And, the voltage on the control grids was several volts positive. Since we had changed all the capacitors, including the coupling capacitors to the control grids, we couldn't come up with an answer as to why this was happening. Several suggestions came from other collectors: a conductive tube (valve) socket that was drawing down the voltages; a high level oscillation in the IF stages; bad tubes (valves) (but we had changed them several times). After several hours of rechecking all the components, we noticed that the filaments on the IF tubes (valves) appeared very bright. A check with the voltmeter revealed that there was almost 7 volts AC present instead of the 4 volts they required.
When replacing the high voltage (EHT) transformer, which also has the filament windings, we had reversed the 6.3 and 4 volt connections. Correcting this solved the problem with the IF stages - apparently the tubes (valves) slowly overheated and starting drawing too much current, thereby driving the plate and screen grid voltages down and causing a positive voltage to appear on the control grids.
Reversing the filament lines also solved the height and width problems, since the 6.3v deflection tubes (valves) were being operated at 4 volts.
A speaker was connected to test the audio, which worked perfectly. The radio also works fine now.
Restoration of this set is now complete.